OpenStudy (anonymous):

how to convert 6.8 meters to feet? ( showing work please)

OpenStudy (akashdeepdeb):

1 feet = 12 inches. 1 inch = 2.54 cm. That means, 1 feet = 12 * 2.54 cm = 30.48 cm That means 1 feet = 30.48 cm = 0.3048 m [Because, 100 cm = 1m] So, if 1 feet = 0.3 m [approximately] Then 1 m = 1/0.3 ft = 10/3 ft = 3.333 ft Therefore, 6.8 m = 6.8 * 3.3333 ft = 22.66 ft. Getting this? :)